An atom is singly ionized so that it acquires a positive charge 1.6 x 10 19 C It then drifts into a device with two parallel plates with a voltage of 175 V across them. This device then accelerates the atom from re 76,000 m/s before it hits the negatively charged plate. What is the mass of the atom? Give your answer in atomic mass units (1.66 x 1027 kg).
Answer #1

Final KE of atom, 0.5mv^2 = qV

0.5*m*76000^2 = 1.6e-19*175

m =  1.6e-19*175*2/76000^2

= 9.695*10^-27 kg

=  9.695*10^-27/1.66e-27 amu

= 5.84 amu